# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

### v1

@@ 5.2. Goal-driven Control Scheme in Deep Phase Manifold
In this section, we demonstrate that the DeepPhase control can be effectively represented as a linear dynamical system within the DeepPhase latent space. Leveraging this characteristic, we propose a goal-driven controller that enables smooth tracking of goal phase feature vectors in a latent space. The Periodic Auto-encoder of the DeepPhase framework converts high-dimensional nonlinear motion in the canonical space to a low-dimensional representation where linear dynamics can be applied, hence suiting learning long-term interactions. Additionally, we introduce a novel feedback loop by utilizing the network’s predictive capability to estimate the goal phase, and establish a relationship between the measurement covariance in the Kalman filter and the bi-directional prediction bias in the spatial domain. This innovative approach significantly enhances guidance in the frequency domain latent, hereby improve the full-body motion coordinate.
@@
(17) $\bm{U}$ $=\left[\begin{array}[]{l}\Delta\mathbf{S}\\ \Delta\mathbf{A}\\ \Delta\mathbf{F}\end{array}\right]=f_{MoE}\left(\mathbf{X}_{t}^{p},\mathbf{X}_{g}^{p},\mathbf{Y}_{t}\right)\sim\bm{C}[\mathbf{X}_{g}^{p},\mathbf{X}_{t}^{p}]^{T}$
where $\beta=0.5$ is a blending weight that enforces the phase update to follow the frequency. Among the state variables, the phase is circularly updated and constrained within the range: $\mathbf{S}_{t}\in[0,1]$, the amplitude $\mathbf{A}_{t}\in[0,1]$, and the frequency $\mathbf{F}_{t}\in[0,\mathbf{F}_{m}]$ where $\mathbf{F}_{m}$ is the Nyquist frequency. Since $|\det(A)|=1$ and the system is fully controllable via $\mathbf{U}$, this system exhibits critical stability and with appropriate control inputs, the system state $\mathbf{X}_{t}^{p}$ can be guided to converge toward the desired goal state $\mathbf{X}_{g}^{p}$ in the latent space. Because the controller network is trained to reduce the difference between the goal phase variables $\mathbf{X}_{g}^{p}$ and the current phase variables $\mathbf{X}_{t}^{p}$, we can approximate its output $\bm{U}$ by a linear transformation of the state error, i.e., $C(\mathbf{X}_{g}^{p}-\mathbf{X}_{t}^{p})$, with $\bm{C}$ as a coefficient matrix. Due to the high-order smoothness of natural movements, the updates $\bm{U}$ are small and bounded.
The control signal $\mathbf{U}$ can be statistically modeled as a zero-mean Gaussian distribution (see Fig. 5). This assumption aligns with the characteristics of human motion sequences, which are often represented by Gaussian distributions (Min and Chai, 2012; Mukai and Kuriyama, 2005). As $U$ follows the Gaussian distribution, the process noise (error from directly inferring the next state by $\mathbf{X}_{t+\Delta t}=\mathbf{A}\mathbf{X}_{t}$) and the measurement noise from the controller network also follows the Gaussian distribution. The approximated linear dynamics of the phase state variables, coupled with the Gaussian noise assumption, allow us to use the Kalman filter to achieve stable goal tracking, as described next.
@@
Ensuring precise and continuous target inputs $X_{g}^{p}$ online is crucial for achieving robust goal tracking and precise convergence to the user’s intended interactions. In our system, as presented in § 5.1, we assume that the goal phase is given. However, using the matched goal provides a biased phase prior due to differences in matching, and can only be provided for a short horizon around the action key-frame. To address this challenge and appropriately lead the character, we develop a framework for accurately estimating goal phases aligned with our planned hand trajectory using the Kalman filter. This filter is optimal for estimating the time-varying goal state under linear system dynamics with Gaussian noise. Our Kalman filter incorporates two measurements, namely predictions from the network ($m=pred$) which is unbiased and well suits time-varying goals, and the matched motion data prior ($m=match$) aligned with the same target action. Algorithm 1 outlines our pipeline.
@@
       $hand$ := (is left-hand task ongoing) ? $lefthand$ : $righthand$
@@
$\sigma_{s}^{2},\sigma_{f}^{2},\sigma_{a}^{2}$ are the process noise of the phase value, frequency, and amplitude, each of which is normalized to 0.2$\times$ of their statistical variance of the entire dataset.
@@
where $Dev(*,*)$ computes the deviation of two 6D configurations and $\mathbf{T}_{j}^{ego}=\left[\bm{r}_{j}^{ego},\bm{t}_{j}^{ego}\right]$, $\mathbf{T}_{j}^{inv}=\left[\bm{r}_{j}^{inv},\bm{t}_{j}^{inv}\right]$ are the ego/goal-centric transformation prediction of the $j$-th key joint; the normalization parameters are set to $\alpha=1/\pi$ and $\beta$ the reciprocal of the body width. Regarding the measurement from the matched motion in the database, the corresponding phase of each frames can be precomputed by the PAE, and the covariance matrix is set to $\mathbf{Q}$.
@@
6. Navigation and Bimanual Planning for Pick and Placing
We now describe how we navigate the character to the target object and then schedule the bimanual manipulation for pick-and-placing. For both tasks, we compute a feature of the environment based on 3D CNN (see § 6.1) and then use it to match motions for navigation (see  § 6.2 and Fig. 2 green sub-system), where a series of locomotion data are stitched to reach the target, and bimanual scheduling (see § 6.3 and Fig. 2, Bimanual Goal Matching block and Goal Coordination block), where a series of keyframes are planned for achieving the pick-and-place actions. Fig. 6 overviews the logic of our controller in the form of a state-machine structure that produces a coordinated fully-body motion.
Figure 6. Overview of our state-machine structure for synthesis coordinate interaction-guiding goals (corresponding to Fig. 2, the bimanual goal coordination before getting the fused goal.). It adaptively allocates coordinated goals for both the hands and the entire body according to the test environment as described in § 6.3. The navigation goal before arrival (see § 6.2) and the interaction goals during manipulation (plot in green capsules) are sequentially generated to guide the DeepPhase Interaction Controller.
Figure 7. Our trajectory planner will detect the cases where the clicking target (teapot) cannot be grasped. To always achieve the interaction task, our system can sequentially remove the obstacle and then re-plan to take the target object out utilizing the available hands, and finally we can place the obstacle back to any clicking position inside the blocks. Among the trajectory visualizations, the object/skeleton from transparent to opaque sequentially shows the in-hand motion; the blue and green trajectories record the approaching and leaving motion, respectively.
@@ 6.1. Environment Features
The goal matching mechanism is to provide a motion prior for grasping/placing an object; to achieve a quick matching, we compute a volumetric scene feature around the goal position by 3D CNN. An autoencoder structure with 3 layers of convolution + max-pooling followed by 3 layers of deconvolution + upscaling is designed. A distance field of the environment pivoted at the object center obtained by fast marching is fed into the network as an input. We extracted the flattened latent code $\bm{z}\in\mathbb{R}^{16}$ together with the hand manipulation state $\bm{c}$ described in Section 4.2 as the matching feature vector from all the goal keyframes in the training set and store them in a KD tree structure.
@@ 6.2. Navigation Matching
We now describe our scheme to guide the character in a current pose to arbitrary goal locations in the scene. Solving such a problem only by motion matching is not feasible especially if the goal location is far away. We cope with this issue by first planning a motion by 2D navigation with breadth-first search, and then finding a sequence of matched motions that can be stitched to reach the goal.
@@
The 2D path from the current position of the character $\bm{p^{\prime}}$ to the goal position and direction $(\bm{p}_{g}^{root},\bm{d}_{g}^{root})$, which can be acquired by leveraging the matched root-to-furniture transformation, is computed by breadth-first search. The cost function for each position $p$ is defined as follows:
(23) $c(\bm{p})=\begin{cases}\infty,&\textit{if}\ \left(\bm{p}\right)in\ obstacles\\ angle(\bm{p}-\bm{p}_{g}^{root},\bm{d}_{g}^{root})/\pi+0.5,&\textit{if}\ ||\bm{p}-\bm{p}_{g}^{root}||_{1}<l\\ 1,&\ \textit{otherwise}\end{cases}$
where $l$ is a threshold set to $0.4m$; the second cost encourages the character to approach the goal from a direction that is parallel to the orientation of the character at the goal. The optimal path is computed using fast marching method.
@@
Given a 2D path computed by the breadth-first search, and the current pose of the character $\bm{p^{\prime}}$, a motion is matched from the dataset to synthesize natural locomotion that follows the path. A motion that has least deviation and has no collision with the surrounding obstacles is selected. As it will be difficult to find a match for a very long path, the path is split $3m$ from the start position and matched for each segment.
For preparing a motion dataset, navigation features defined by $\bm{N}=\{\bm{d}_{g}^{root},\bm{p}_{g}^{root},\bm{r}^{left},\bm{p}^{left},\bm{r}^{right},\bm{p}^{right},\bm{r}^{hip},\bm{p}^{hip},\bm{v}^{root}\}$ are computed for every motion between adjacent keyframes in the database, where $\bm{d}_{g}^{root}\in\mathbb{R}^{2}$ and $\bm{p}_{g}^{root}\in\mathbb{R}^{2}$ are 2D orientation/position at the goal and $\bm{r}^{keyJ}\in\mathbb{R}^{6}$, $\bm{p}^{keyJ}\in\mathbb{R}^{3}$ $\bm{v}^{root}\in\mathbb{R}^{2}$ are key joint orientation/position and root velocity relative to the root coordinate at the beginning of each motion clip. These features are used as the key to search for motion clips during navigation.
@@ 6.3. Bimanual Task Scheduler
Humans use both hands to achieve pick-and-place tasks efficiently in living environments, e.g. using one hand to open the door of the cabinet while taking out an object; to produce such movements, we propose a bimanual task scheduler that schedules the optimal series of actions by the full body.
@@
1. Detecting the target and auxiliary objects: The user clicks the goal object with the mouse, and then the system detects the target object (denoted as $o_{g}$) and all the other objects involved, which we call auxiliary objects $\{o_{d}\}$, such as the obstacles, and the nearby container (whose manipulable part like a door/drawer is also denoted as $o_{d}$) by ray-casting.
The system then checks if any auxiliary object is required to be manipulated. In addition to handling the closed cabinets, Fig. 7 shows another no-solution case to directly take out the inner teapot, where our system automatically removes the obstacle object before taking the target one. This is determined by whether the target object can be taken outside the cabinet without collision, based on the planned trajectory.
@@
The scheduling is done according to the availability of the hands: If both hands are free, the goal prior matching algorithm described in Section 6.1 is conducted separately for $o_{g}$ and $o_{d}$, resulting in 4 pairs of hand-object interaction matches denoted by $\{M(H_{\{l,r\}},o_{g})$, $M(H_{\{l,r\}},o_{d})\}$ where $\{l,r\}$ represents the left and right hand. If only one hand is free, the opening and grasping tasks are allocated sequentially to the free hand; if two hands are free, the total goal costs are computed for all potential schedules, including sequentially opening and grasping with either hand and co-temporally opening and grasping with two hands. The keyframes to grasp the auxiliary object and the target object are matched such that they minimize the designed deviation cost $C$ for all possible combinations of keyframes for grasping the auxiliary object and the target object:
@@
3. Runtime Motion Synthesis: Given the series of keyframes for grasping/releasing extracted from the dataset, they are adjusted to the runtime scene by IK and used for motion synthesis. They are given to the trajectory planner to produce the trajectories of the hands and are also used as the goal poses for the DeepPhase controller. We prepare a state machine for these procedures including the 2D path planner to approach the goal as shown in Fig. 6.
@@ 7. Data Capture and Labeling
Figure 8. Snapshots of our CHOICE dataset, which incorporates different tasks like (a) cooking with ingredients and seasoning, (b) making milk tea and pouring to multiple mugs, (c) cleaning the table, and (d) other bimanual manipulations that co-temporally manipulating two objects.
Table 2. The distribution of action labels and task descriptions for the key joints. ”Cook” refers to the motion operating the pan/spatula, and ”Pour” includes both pouring drinking and the liquid ingredients.
Key Joint Action Task Seconds %
Grasp/Place 3451 12.7
Object Manipulate Cook 2845 10.5
Drink 98 0.4
Pour 3075 11.3
Door Manipulate Door opening/closing 3046 11.2
Draw Drawer Drag 1731 6.4
Idle Turn/Stand/Squat 23352 85.9
Walk Approach 3845 14.1
@@
We extensively captured and post-processed 150 motion sequences using the Vicon Shógun  [^1] https://www.vicon.com/software/shogun/ motion capture system. The dataset comprises long-term, delicate interactions in lifelike scenes, such as a corridor with a drinking bar (Fig. 8 (d)), and a well-equipped kitchen room (Fig. 8 (a)). We approximate various involved furniture by configuring up iron frames, the plains in drawers and tables are formed by wired mesh, and all the articulated doors are glass-made. This setup ensured that every contained object was visible to the surrounding cameras. After filtering the initial data, we post-process our captures in Vicon Shógun by minimizing finger-object penetration and ensuring contact during grasping. We fit an SMPL (Loper et al., 2023) body model to the subject using SOMA (Ghorbani and Black, 2021), and export it to the Vicon Shógun body template with a hand skeleton.
@@
Each motion sequence in our captured dataset achieves a high-level semantic goal, e.g., cooking a seasoned steak, making milk tea and drinking, or cleaning a crowded table, with an average duration of 3 minutes. The dataset exhibits substantial diversity in terms of hand movements, object types (including 14 pieces of heavy furniture, 12 freely movable objects, 6 doors, and 4 drawers) and spatial layouts, which vary across cabinets and drawers at different widths and heights, object-object relations, obstacles, and placement heights ranged from 20cm to 200cm.
@@
In addition to the sequence-level labelling, we annotate each frame of the three key joints with action labels. The hip is labelled as either ”idle” or ”walk”, and each hand is labelled with actions like ”open a door”, ”open a drawer”, or ”approaching an object”, along with a contact label with the object. A combination of labels for the three key-joints effectively describes the current action type (see Table 2).
@@
We further utilize the PAE’s encoder to extract the phase latent vector that implicitly represents the full-body motion information of the $[-1s,1s]$ time window pivoted on each frame. Consequently, each frame is described by its current state and the next goal state, with each state consisting of its phase latent vector, action labels, and the three key-joint transformations.

### v2

@@ 5.2. Goal-driven Control Scheme in Deep Phase Manifold
In this section, we demonstrate that the DeepPhase control can be effectively represented as a linear dynamical system within the DeepPhase latent space. Leveraging this characteristic, we propose a goal-driven controller that enables smooth tracking of goal phase feature vectors in a latent space. The Periodic Auto-encoder of the DeepPhase framework converts high-dimensional nonlinear motion in the canonical space to a low-dimensional representation where linear dynamics can be applied, hence suiting learning long-term interactions. Additionally, we introduce a novel feedback loop by utilizing the network’s predictive capability to estimate the goal phase, and establish a relationship between the measurement covariance provided to the Kalman filter and the bi-directional prediction bias in the spatial domain. This innovative approach significantly enhances guidance in the frequency domain latent, thereby improving the full-body motion coordination.
@@
(17) $\mathbf{U}$ $=\left[\begin{array}[]{l}\Delta\mathbf{S}\\ \Delta\mathbf{A}\\ \Delta\mathbf{F}\end{array}\right]=f_{MoE}\left(\mathbf{X}_{t}^{p},\mathbf{X}_{g}^{p},\mathbf{Y}_{t}\right)\sim\bm{C}[\mathbf{X}_{g}^{p},\mathbf{X}_{t}^{p}]^{T}$
where $\beta=0.5$ is a blending weight that enforces the phase update to follow the frequency. Among the state variables, the phase features are circularly updated and constrained within the range: the phase value $\mathbf{S}_{t}\in[0,1]$, the amplitude $\mathbf{A}_{t}\in[0,1]$, and the frequency $\mathbf{F}_{t}\in[0,\mathbf{F}_{m}]$, where $\mathbf{F}_{m}$ is the Nyquist frequency. Since $|\det(A)|=1$ and the system is fully controllable via $\mathbf{U}$, this system exhibits critical stability, and with appropriate control inputs, the system state $\mathbf{X}_{t}^{p}$ can be guided to converge toward the desired goal state $\mathbf{X}_{g}^{p}$ in the latent space. Because the controller network is trained to reduce the difference between the goal phase variables $\mathbf{X}_{g}^{p}$ and the current phase variables $\mathbf{X}_{t}^{p}$, we can approximate its output of phase update $\mathbf{U}$ by a linear transformation of the state error, i.e., $\mathbf{C}(\mathbf{X}_{g}^{p}-\mathbf{X}_{t}^{p})$, with $\mathbf{C}$ as a coefficient matrix. Due to the high-order smoothness of natural movements, the updates $\mathbf{U}$ are small and bounded.
The control signal $\mathbf{U}$ can be statistically modeled as a zero-mean Gaussian distribution (see Fig. 5). This assumption aligns with the characteristics of human motion sequences, which are often represented by Gaussian distributions (Min and Chai, 2012; Mukai and Kuriyama, 2005). As $U$ follows the Gaussian distribution, the process noise (error from directly inferring the next state by $\mathbf{X}_{t+\Delta t}=\mathbf{A}\mathbf{X}_{t}$) and the measurement noise from the controller network also follow the Gaussian distribution. The approximated linear dynamics of the phase state variables, coupled with the Gaussian noise assumption, allow us to use the Kalman filter to achieve stable goal tracking, as described next.
@@
Ensuring precise and continuous target inputs $\mathbf{X}_{g}^{p}$ online is crucial for achieving robust goal tracking and precise convergence to the user’s intended interactions. In our system, as presented in § 5.1, we assume that the goal phase is given. However, using the matched goal provides a biased phase prior due to differences in matching, and can only be provided for a short horizon around the action keyframe. To address this challenge and appropriately lead the character, we develop a framework for accurately estimating goal phases aligned with our planned hand trajectory using the Kalman filter. This filter is optimal for estimating the time-varying goal state under linear system dynamics with Gaussian noise. Our Kalman filter incorporates two measurements, namely predictions from the network ($m=pred$) which is unbiased and suits time-varying goals well, and the matched motion data prior ($m=match$) aligned with the same target action. Algorithm 1 outlines our pipeline.
@@
 handhand := (is left-hand task ongoing) ? lefthandlefthand : righthandrighthand
@@
[…]
@@
where Dev(,)Dev(*,*) computes the deviation of two 6D configurations and $\mathbf{T}_{j}^{ego}=\left[\bm{r}_{j}^{ego},\bm{t}_{j}^{ego}\right]$, $\mathbf{T}_{j}^{inv}=\left[\bm{r}_{j}^{inv},\bm{t}_{j}^{inv}\right]$ are the ego/goal-centric transformation prediction of the $j$-th key joint; the normalization parameters are set to $\alpha=1/\pi$ and $\beta$ the reciprocal of the body width. Regarding the measurement from the matched motion in the database, the corresponding phase of each frame can be precomputed by the PAE, and the covariance matrix is set to $\mathbf{Q}$.
@@
6. Navigation and Bimanual Goal Scheduling
We now describe how we navigate the character to the target object and then schedule the bimanual interaction goals for pick-and-placing. This procedure can be decomposed to a goal matching subsystem (see Fig. 2 green sub-system) and a state machine (see Fig. 2 grey block of Goal Coordination). Before running the goal matching, we first compute a feature of the environment based on 3D CNN (see § 6.1). Next, we use this feature to retrieve and stitch a series of locomotion snippets, driving the navigation to the goal object (see  § 6.2 and Fig. 2 cyan block). Once the character arrives, the bimanual goal matching module (see § 6.3 and Fig. 2, green block) plans a sequence of keyframes to execute the pick-and-place actions. Fig. 6 overviews the logic of our goal coordination module in the form of a state-machine structure that produces coordinated full-body motion.
Figure 6. Overview of our state-machine structure for synthesis coordinate interaction-guiding goals (corresponds to Fig. 2, the goal coordination block before getting the fused goal). It adaptively allocates coordinated goals for both the hands and the entire body according to the test environment as described in § 6.3. The navigation goal before arrival (see § 6.2) and the interaction goals during manipulation (plot in green capsules) are sequentially generated to guide the DeepPhase interaction controller.
Figure 7. Our trajectory planner will detect the cases where the clicking target (teapot) cannot be grasped. To always achieve the interaction task, our system can sequentially remove the obstacle and then re-plan to take the target object out utilizing the available hands, and finally, we can place the obstacle back to any clicking position inside the blocks. Among the trajectory visualizations, the object/skeleton from transparent to opaque sequentially shows the in-hand motion; the blue and green trajectories record the approaching and leaving motion, respectively.
@@ 6.1. Environment Features
The goal matching mechanism is to provide a motion prior for grasping/placing an object; to achieve a quick matching, we compute a volumetric scene feature around the goal position by 3D CNN. An auto-encoder structure with 3 layers of convolution + max-pooling, followed by 3 layers of deconvolution + upscaling, is designed. A distance field of the environment, pivoted at the object center and computed by fast marching, is fed into the network as an input. We extracted the flattened latent code $\bm{z}\in\mathbb{R}^{16}$ together with the hand manipulation state $\bm{c}$ described in Section 4.2 as the matching feature vector from all the goal keyframes in the training set and store them in a KD tree structure.
@@ 6.2. Navigation Matching
We now describe our scheme to guide the character in a current pose to arbitrary goal locations in the scene. Solving such a problem only by motion matching is not feasible, especially if the goal location is far away. We cope with this issue by first planning a motion by 2D navigation with breadth-first search, and then finding a sequence of matched motions that can be stitched to reach the goal.
@@
The 2D path from the current position of the character $\bm{p^{\prime}}$ to the goal position and direction $(\bm{p}_{g}^{root},\bm{d}_{g}^{root})$, which can be acquired from the matched destination keyframe with hand-object operation, is computed by a breadth-first search. The cost function for each position $p$ is defined as follows:
[…]
where $l$ is a threshold set to $0.4m$; the second cost encourages the character to reach the goal from a direction that is parallel to the facing direction of the character at the goal. The optimal path is computed using the fast marching method.
@@
Given a 2D path computed by breadth-first search and the current pose of the character $\bm{p^{\prime}}$, goal waypoints of the three keyjoints are matched from the dataset to synthesize natural locomotion that follows the path. As it will be difficult to find a match for a very long path, we track the path in segments by querying a series of goal positions uniformly sampled within each $3m$ segment from the start root position. In each segment, the keyjoint goal with the least deviation and without any collision with the surrounding obstacles is then selected.
For preparing a motion dataset, navigation features defined by $\bm{N}=\{\bm{d}_{g}^{root},\bm{p}_{g}^{root},\bm{r}^{left},\bm{p}^{left},\bm{r}^{right},\bm{p}^{right},\bm{r}^{hip},\bm{p}^{hip},\bm{v}^{root}\}$ are computed for every motion between adjacent keyframes in the database, where $\bm{d}_{g}^{root}\in\mathbb{R}^{2}$ and $\bm{p}_{g}^{root}\in\mathbb{R}^{2}$ are 2D orientation/position at the goal and $\bm{r}^{keyJ}\in\mathbb{R}^{6}$, $\bm{p}^{keyJ}\in\mathbb{R}^{3}$ , $\bm{v}^{root}\in\mathbb{R}^{2}$ are keyjoint orientation/position and root velocity relative to the root coordinate at the beginning of each motion clip. These features are used as the key to search for motion clips during navigation.
@@ 6.3. Bimanual Task Scheduler
Humans use both hands to achieve pick-and-place tasks efficiently in living environments, e.g., using one hand to open the door of the cabinet while taking out an object with the other hand; to produce such movements, we propose a bimanual task scheduler that schedules the optimal series of actions by the full body.
@@
1. Detecting the target and auxiliary objects: The user clicks the target object with the mouse, and then the system detects the target object (denoted as $o_{g}$) and all the other objects involved, which we call auxiliary objects $\{o_{d}\}$, such as the obstacles, and the container that contains the target object (whose manipulable part like a door/drawer is also denoted as $o_{d}$) by ray-casting.
The system then checks if any auxiliary object needs to be manipulated. In addition to handling the closed cabinets, Fig. 7 shows another no-solution case to directly take out the inner teapot, where our system automatically controls the character to remove the obstacle object before taking the target one. This is determined by whether the target object can be taken outside the cabinet without collision, based on the planned trajectory.
@@
The task scheduling is done according to the availability of the hands: If both hands are free, the goal prior matching algorithm described in Section 6.1 is conducted separately for $o_{g}$ and $o_{d}$, resulting in 4 pairs of hand-object interaction matches denoted by $\{M(H_{\{l,r\}},o_{g})$, $M(H_{\{l,r\}},o_{d})\}$ where $\{l,r\}$ represents the left and right hand. If only one hand is free, the opening and grasping tasks are allocated sequentially to the free hand.
If two hands are free, the total goal costs are then computed for all potential schedules, including sequentially opening and grasping with either hand and co-temporally opening and grasping with two hands. The keyframes to grasp the auxiliary object and the target object are matched such that they minimize the designed deviation cost $C$ for all possible combinations of keyframes for grasping the auxiliary object and the target object:
@@
3. Runtime Motion Goal Synthesis: Given a sequence of grasping/releasing keyframes extracted from the dataset, we adjust their corresponding keyjoint transformations to the runtime scene after a collision-solving process that pushes back any penetrations of the hip or the idle hand along the penetration normal. The adjusted goal on the hand-object contact keyframe is then passed to the trajectory planner to generate the hand trajectories, which we compose with the other goal keyjoints to drive the DeepPhase interaction controller. We finally establish a state machine for these procedures, including the navigation stage before scheduling the manipulation, as shown in Fig. 6.
@@ 7. Data Capture and Labeling
Figure 8. Snapshots of our CHOICE dataset, which incorporates different tasks like (a) cooking with ingredients and seasoning, (b) making milk tea and pouring into multiple mugs, (c) cleaning the table, and (d) other bimanual manipulations that co-temporally operate two objects. In (e), we visualize the pose variations in datasets after T-SNE clustering.
Table 2. We list the distribution of action labels and task descriptions for the key joints of our dataset, and evaluate the motion diversity and smoothness by class. ”Cook” denotes the motion of operating the pan/spatula, and ”Pour” encompasses both pouring drinking and the liquid ingredients. The ”Div.” metric evaluates the motion diversity by Average Pairwise Distance (APD) in $cm$. We sample and compute this metric consistently with HuMoR (Rempe et al., 2021), where its motion latent reconstructs a 95 APD from the huge AMASS dataset (Mahmood et al., 2019). We also measure the motion smoothness by reporting the mean acceleration (in $cm/s^{2}$) and jerk (in $m/s^{3}$) averaged among all non-finger joints.
Action Task Sec. % Div. Acc. Jerk
Grasp/Place 3451 12.7 96.2 90.7 62.1
Object Cook 2845 10.5 91.1 76.4 57.1
Drink 98 0.4 88.1 89.1 50.7
Pour 3075 11.3 74.8 86.1 49.9
Door Open/Close 3046 11.2 100.3 87.4 50.9
Draw Drag/Push 1731 6.4 130.5 80.6 50.5
Idle Stand/Squat 23352 85.9 95.8 92.9 55.5
Walk Approach/Leave 3845 14.1 84.2 146.2 62.4
Table 3. Dataset evaluation on grasping precision. Our evaluation is computed by averaging the maximum penetration volume and depth, and floating displacement during each pick-and-place.
p.vol. $(cm^{3})$ p.dist. $(cm)$ disp. $(cm)$
1.74 0.22 0.61
@@
We extensively captured and post-processed 150 motion sequences using the Vicon Shógun  [^1] https://www.vicon.com/software/shogun/ motion capture system. The dataset comprises long-term, delicate interactions in lifelike scenes, such as a corridor with a drinking bar (Fig. 8(d)), and a well-equipped kitchen room (Fig. 8(a)). We approximate various involved furniture by configuring up iron frames, the surfaces in drawers and tables are formed by wired mesh, and all the articulated doors are made of glass. This setup ensures that every contained object is visible to the surrounding cameras. After filtering the initial data, we post-process our captures in Vicon Shógun by minimizing finger-object penetration and ensuring contact during grasping. We fit an SMPL (Loper et al., 2023) body model to the subject using SOMA (Ghorbani and Black, 2021), and export it to the Vicon Shógun body template with a hand skeleton.
The breakdown of our dataset is shown in Tab. 2: Following the detailed MoCap procedure described in our supplementary, our dataset achieves high smoothness, as detailed in Tab. 2, along with precise manipulation, as shown in Tab. 3.
@@
Each motion sequence in our captured dataset achieves a high-level semantic goal, e.g., cooking a seasoned steak, making milk tea and drinking, or cleaning a crowded table, with an average duration of 3 minutes. The dataset exhibits substantial diversity in terms of hand movements, object types (including 14 pieces of heavy furniture, 12 freely movable objects, 6 doors, and 4 drawers), and spatial layouts, which vary across cabinets and drawers at different widths and heights, object-object relations, obstacles, and placement heights ranging from 20cmcm to 200cmcm. Our dataset offers pose diversity comparable to the OMOMO (Li et al., 2023) dataset (see Fig. 8(e)), and captures rich motion variations for each action category, as quantified in the ”Div” column of  Tab. 2.
@@
In addition to the sequence-level labelling, we annotate each frame of the three key joints with action labels. The hip is labelled as either ”idle” or ”walk”, and each hand is labelled with actions like ”open a door”, ”open a drawer”, or ”approach an object”, along with a contact label with the object. A combination of labels for the three keyjoints effectively describes the current action type (see Tab. 2).
@@
We further utilize the PAE’s encoder to extract the phase latent vector that implicitly represents the full-body motion information of the [1s,1s][-1s,1s] time window pivoted on each frame. Consequently, each frame is described by its current state and the next goal state, with each state consisting of its phase latent vector, action labels, and the three key-joint transformations.